OpenStudy (anonymous):

@Data_LG2 Hi Could you please help me with a couple of Physics Problems? You were very helpful last time

OpenStudy (anonymous):

\[\color{green}{T = 2\pi \sqrt{\dfrac{l}{g}}}\] For #1 solve for g.

OpenStudy (anonymous):

Yes, that looks fine. Well done.